The perimeter of a rectangle is 144 ft. The ratio of the width to the length is 3:5. Find the dimensions (length and width) of the rectangle.
Answer by jim_thompson5910(35256)

"The ratio of the width to the length is 3:5"
Width: 3x
Length: 5x
for some number x
-------------------------------------------------------

2*(L+W) = P

2*(3x+5x) = 144

2*8x = 144

16x = 144

x = 144/16

x = 9

-------------------------------------------------------

Width: 3x = 3*9 = 27
Length: 5x = 5*9 = 45

The length is 45, the width is 27.